# Weakly Supervised Veracity Classification with LLM-Predicted Credibility Signals

### v2

@@ 1 Introduction
To address these issues, prior work has employed weakly supervised methods that leverage indirect learning signals to classify the veracity of content without relying on annotated data. Current methods use weak signals as a combination of simple syntactic features (e.g., count of words) and user engagement with the misinformation content (e.g., number of shares) [7, 8, 9]. The latter is particularly ineffective, as models that depend on engagement features require the content to be spread and interacted with before the model can accurately detect its deceptive nature, by which time the misinformation narrative has already caused harm. In spite of the simplicity of the aforementioned signals, the challenge of integrating more sophisticated signals poses a paradox: complex signals demand specialised models and annotated datasets for accurate extraction [10], which undermines the premise of employing weak supervision in the first place.
@@
Our contribution with this work is the proposal of Prompted weAk superviSion wiTh crEdibility signaLs (Pastel), an approach modelled on the verification process typically adopted by journalists and fact-checkers, who assess the veracity of online content using a wide range of credibility signals [^2] See the report by W3C-CWCG [14] for an overview of credibility signals.. We leverage the task-agnostic capabilities of LLMs to extract nineteen sophisticated credibility signals from news articles in a zero-shot setting (i.e., without training the model with ground truth labels). These signals are then aggregated into a binary (misinformation/non-misinformation) veracity label using weak supervision.
@@
The rest of the paper is structured as follows: Section 2 presents an overview of relevant previous work. Section 3 describes our proposed method. The experimental setup is presented in Section 4, whilst results are discussed in Section 5. In Section 6, we analyse and discuss the predicted credibility signals. Section 7 presents a discussion of implications of this work, points to future work, and makes concluding remarks. We make our code and data fully available [15].
@@ 2.2 Credibility Signals
The term credibility signals refers to a wide range of measurable heuristics that collectively help journalists assess the overall trustworthiness of information. Examples of credibility signals include the analysis of article titles [35], writing style [36], rhetorical structure [37], linguistic features [38], emotional language [39], biases [40], and logical fallacies and inferences [41]. Additionally, credibility signals comprise meta-information that extends beyond the textual content of the article, such as the author’s reputation and external references [42].
@@ 3 Prompted Weak Supervision with Credibility Signals (Pastel)
Figure 1: Prompted weak supervision with credibility signals (Pastel).
@@ 3.1 Credibility Signals Considered
We leverage nineteen credibility signals, all of which have been shown to be relevant for assessing content veracity. Table 1 displays the nineteen signals we considered, providing a solid foundation of well-defined and validated indicators of content credibility. Note that all our signals are formulated so that their presence in the content indicates a lack of credibility.
@@
The vast majority of the signals used in our experiments were proposed by the W3C (the Web Standards Organisation) Credible Web Community Group [14], who defined numerous credibility indicators to help users and machines identify trustworthy content, i.e. content that is reliable, accurate, and shared in good faith. The aim of our work is not to propose new credibility signals but to use those already established by subject matter experts and demonstrate how they can enhance automatic veracity classification.
@@ 5.1 In-domain Classification
In this scenario, models are trained and evaluated with in-domain data, i.e., the train and test sets are derived from the same dataset. Table 3 presents the classification results for the proposed baselines.
@@ 5.2 Cross-domain Classification
Although Pastel consistently outperforms RoBERTa, the difference is less pronounced for datasets within the same domain, particularly entertainment news. For instance, when RoBERTa is trained on GossipCop and tested on Celebrity, it achieves an F1macro score of $0.74$, which is $0.07$ behind Pastel. When trained on Celebrity and evaluated on GossipCop, both models score $0.69$. In the political domain, the performance gap is more significant. Training on PolitiFact and evaluating on FakeNewsAMT results in an F1macro score of $0.54$ for RoBERTa, $0.22$ lower than Pastel. Similarly, training on FakeNewsAMT and testing on PolitiFact yields a score of $0.40$ for RoBERTa, which is $0.27$ below Pastel.
@@ 5.3 Error Analysis
Table 6: Relative frequency of credibility signals triggered in TP and FN predictions. Percent decrease indicated within parenthesis.
Credibility Signals TP FN
[…]
The statistics indicate that all $19$ signals occur less frequently in FN predictions compared to TP. On average, $7.6$ credibility signals are triggered in TP predictions, whereas only $0.9$ signals are triggered in FN predictions, representing a significant decrease of $87.6\%$. A reduction of more than $70\%$ in frequency is observed for $16$ signals, while Call to Action, Inference, and Reported by Other Sources show smaller decreases of $39.1\%$, $40.3\%$, and $51.2\%$, respectively.
@@ Acknowledgements
We thank Freddy Heppel, Ivan Srba, and Ben Wu for their valuable feedback.

### v3

@@ 1 Introduction
To address these issues, prior work has employed weakly supervised methods that leverage indirect learning signals to classify the veracity of content without relying on annotated data. Current methods use weak signals as a combination of simple syntactic features (e.g., count of words) and user engagement with the misinformation content (e.g., number of shares) [7, 8, 9]. The latter is particularly ineffective, as models that depend on engagement features require the content to be spread and interacted with before the model can accurately detect its deceptive nature, by which time the misinformation narrative has already caused harm. In spite of the simplicity of the aforementioned signals, the challenge of integrating more sophisticated signals (e.g. credibility signals [^2] See the report by W3C-CWCG [10] for an overview of credibility signals. defined by experts) poses a paradox: complex signals demand specialised models and annotated datasets for accurate extraction [11], which undermines the premise of employing weak supervision in the first place.
@@
Our contribution with this work is the proposal of Prompted weAk superviSion wiTh crEdibility signaLs (Pastel), an approach modelled on the verification process typically adopted by journalists and fact-checkers, who assess the veracity of online content using a wide range of credibility signals. We leverage the task-agnostic capabilities of LLMs to extract nineteen sophisticated credibility signals from news articles in a zero-shot setting (i.e., without training the model with ground truth labels). These signals are then aggregated into a binary (misinformation/non-misinformation) veracity label using weak supervision.
@@
The remainder of this paper is structured as follows: Section 2 presents an overview of relevant previous work. Section 3 describes our proposed method. The experimental setup is presented in Section 4, whilst results are discussed in Section 5. In Section 6, we analyse and discuss the predicted credibility signals. Section 7 presents a discussion of implications of this work, points to future work, and makes concluding remarks. We make our code and data fully available [15].
@@ 2.2 Credibility Signals
The term credibility signals refers to a wide range of measurable heuristics that collectively help journalists assess the overall trustworthiness of information. Examples of credibility signals include the analysis of article titles [35], writing style [36], rhetorical structure [37, 38], linguistic features [39], emotional language [40], biases [41], and logical fallacies and inferences [42]. Additionally, credibility signals comprise meta-information that extends beyond the textual content of the article, such as the author’s reputation and external references [43].
@@ 3 Prompted weAk Supervision wiTh crEdibility signaLs (PASTEL)
Figure 1: Illustration of PASTEL.
@@ 3.1 Credibility Signals Considered
We leverage nineteen credibility signals, all of which have been shown to be relevant for assessing content veracity. Table 1 displays these signals, that provide a solid foundation of well-defined and validated indicators of content credibility. Note that all our signals are formulated so that their presence in the content indicates a lack of credibility.
@@
The vast majority of the signals used in our experiments were proposed by the W3C (the Web Standards Organisation) Credible Web Community Group [10], who defined numerous credibility indicators to help users and machines identify trustworthy content, i.e. content that is reliable, accurate, and shared in good faith (see Table 1). The aim of our work is not to propose new credibility signals but to use those already established by subject matter experts and demonstrate how they can enhance automatic veracity classification.
@@ 5.1 In-domain Classification
In the in-domain scenario, models are trained and evaluated with in-domain data, i.e., the train and test sets are derived from the same dataset. Table 3 presents the classification results for the proposed baselines.
@@ 5.2 Cross-domain Classification
Although Pastel consistently outperforms RoBERTa, the difference is less pronounced for datasets within the same domain, particularly entertainment news. For instance, when RoBERTa is trained on GossipCop and tested on Celebrity, it achieves an F1macro score of $0.74$, which is $0.07$ lower than Pastel. When trained on Celebrity and evaluated on GossipCop, both models score $0.69$. In the political domain, the performance gap is more significant. Training on PolitiFact and evaluating on FakeNewsAMT results in an F1macro score of $0.54$ for RoBERTa, $0.22$ lower than Pastel. Similarly, training on FakeNewsAMT and testing on PolitiFact yields a score of $0.40$ for RoBERTa, which is $0.27$ below Pastel.
@@ 5.3 Error Analysis
Table 6: Relative frequency of credibility signals triggered in True Positive (TP) and False Negative (FN) predictions. Percent decrease indicated within parenthesis.
Credibility Signal TPs FNs
Emotional Valence 0.52 0.00(↓100%)
Clickbait 0.29 0.00(↓100%)
Expert Citation 0.55 0.00(↓100%)
Evidence 0.56 0.00(↓100%)
Source Credibility 0.25 0.00(↓100%)
Bias 0.39 0.01(↓97.4%)
Document Citation 0.73 0.02(↓97.3%)
Incivility 0.26 0.01(↓96.2%)
Sensationalism 0.69 0.03(↓95.7%)
Polarising Language 0.39 0.02(↓94.9%)
Misleading about content 0.57 0.05(↓91.2%)
Explicitly Unverified Claims 0.32 0.03(↓90.6%)
Incorrect Spelling 0.19 0.02(↓89.5%)
Impoliteness 0.08 0.01(↓87.5%)
Informal Tone 0.34 0.08(↓76.5%)
Personal Perspective 0.38 0.09(↓76.3%)
Reported by Other Sources 0.78 0.38(↓51.3%)
Call to Action 0.04 0.02(↓50.0%)
Inference 0.24 0.14(↓41.7%)
Total 7.57 0.91(↓88.0%)
The statistics indicate that all $19$ signals occur less frequently in FN predictions compared to TP. On average, $7.57$ credibility signals are triggered in TP predictions, whereas only $0.91$ signals are triggered in FN predictions, representing a significant decrease of $88.0\%$. A reduction of more than $70\%$ in frequency is observed for $16$ signals, while Reported by Other Sources, Call to Action, and Inference show smaller decreases of $51.3\%$, $50.0\%$, and $41.7\%$, respectively.
@@ Acknowledgements
We thank Freddy Heppel, Ivan Srba, and Ben Wu for their valuable feedback. We also acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing.